OpenStudy (bekkah323):

A multiple choice test contains 12 questions, each with four possible answers. Assume that a student randomly guesses on each question. Let the random variable x equal the number of questions that are answered correctly.Generate the binomial probability distribution for x. How do you figure out the success probability?

OpenStudy (bekkah323):

If there are four possible answers to each question, what is the probability that you guess correctly, just on one question?

OpenStudy (kropot72):

The probability that you guess correctly on exactly one question is found from: \[\large P(1\ out\ of\ 12\ correct)=12C1\times\frac{1}{4}\times(\frac{3}{4})^{11}\]

OpenStudy (bekkah323):

That is the formula for calculating probabilities, what I was wondering is if the trial size in this question is 12 and the probability for success was 0.25

OpenStudy (princeharryyy):

P(x correct out of 12) = 12CX *(p)^x *(1-p)^(12-x) where small p = 1/12

OpenStudy (princeharryyy):

Well sorry. Small p has to be 1/4 as probability of getting 1 question right is 0.25. All the rest binomial distribution remains same.

OpenStudy (princeharryyy):

@bekkah323